OpenStudy (anonymous):

Otto mixed spices to create his own Italian seasoning. He used %10 garlic powder. Which of the following shows the percent equation and solution for how many tablespoons of garlic powder he uses to create 20 tablespoons of the seasoning?

OpenStudy (nottim):

you should show what is following

OpenStudy (anonymous):

A.) 20 = 0.1 * a; a= 2 B.) A = 0.1 * 20; a = 200 C.) 20 = 0.1 * a; a = 200 D.) A = 0.1 * 20; a = 2

OpenStudy (nottim):

ok. i think i understand. first off, what is your opinion jade? Its best if we work off wat you nkno first

OpenStudy (anonymous):

I know something like this but it's not working when I try it on this problem.

OpenStudy (nottim):

ok.well, if we want 20 tablespoons of something, and we want to figure out how much garlic powder is in it (where its always 10), how do we figure it out? Really, how do we get 10% of a certain value?

OpenStudy (anonymous):

Uhh, I'm not really sure.

OpenStudy (nottim):

ok. here's anotehr example. you know taxes? how do you get 10% of $4?

OpenStudy (anonymous):

Don't you have to turn the % into a decimal?

OpenStudy (nottim):

yup

OpenStudy (anonymous):

And then multiply it? I'm not sure what you do after the decimal

OpenStudy (nottim):

yes, yo ugot it.

OpenStudy (anonymous):

Really?!? :D

OpenStudy (anonymous):

What would you do in the problem to solve it? Would you do the same thing?

OpenStudy (nottim):

yes. 0.10*20

OpenStudy (anonymous):

Okay thank you! :D

OpenStudy (nottim):

yeah. hey, could you give me a virtual shiny? their really all i have to live for

OpenStudy (anonymous):

How? I just joined.

OpenStudy (nottim):

ya see the good answer button to the right?

OpenStudy (anonymous):

Oh okay, and yes I can :D

OpenStudy (nottim):

yayy. thank you. nottim away